OpenStudy (anonymous):

Find the volume of an 8 in. tall right triangular pyramid with a base hypotenuse of 5 in. and base leg of 3 in. PLEASE HELP!

OpenStudy (anonymous):

1/6Bh where B is the area of the base

OpenStudy (anonymous):

its very simple. all you do is you multiply 8 times 5 times 3 and your answer will be 120

OpenStudy (anonymous):

That answer isn't right

OpenStudy (anonymous):

well, it said the correct answer is 16in cubed, but i don't understand how.

OpenStudy (anonymous):

Okay, to first you need the area of the base of the triangle. If the base of the triangle has a hypotenuse of 5 in and a leg of 3, you need to use the Pythagorean therom to find the length of the other leg of the base so 5^2=3^2+x^2, solve for x. Then use 1/2b*h to find the area of base, divide that by 6 and multiply by the height of the actual pyrmid

OpenStudy (anonymous):

Okay, to first you need the area of the base of the triangle. If the base of the triangle has a hypotenuse of 5 in and a leg of 3, you need to use the Pythagorean therom to find the length of the other leg of the base so 5^2=3^2+x^2, solve for x. Then use 1/2b*h to find the area of base, divide that by 6 and multiply by the height of the actual pyrmid

OpenStudy (anonymous):

oh well you 1/2 the width and multiply the width by the base and the height

OpenStudy (anonymous):

Did you get that? I feel like these other two are trying to give you a hard time

OpenStudy (anonymous):

who

OpenStudy (anonymous):

no, i got it, thank you so much brycoope!

OpenStudy (anonymous):

No problem man

OpenStudy (anonymous):

no problem man

OpenStudy (anonymous):

thank you cshalvey. but, could we please take this off of my question? thank you!(: